A car travels 400 miles on level terrain in the same amount of time it travels 160 miles on mountainous terrain. If the rate of the car is 30 miles per hour less in the mountains than on level ground, find its rate in the mountains.
Answer by checkley75(3666)

distance=rate*time or time=distance/rate
seeing as the times are the same we have the following equatyion:
400/r=160/(r-30) cross multiplying we get:
160r=400r-1200
-240r=-1200
r=-1200/-24
r=50 mph on level ground & 50-30=20 mph on the mountain.
proof
400/50=160/20
8=8